solve a number ifs halved and the outcome then increased by nine. The result is 63. What is the original number?
Answer by saltzy57(8)

Let X = the number we trying to find, to halve a number we divide by 2 and then we add 9 the result = 63.
x/2 + 9 = 63
x/2 = 63-9
x/2 = 54 multiply both sides by 2
x = 108
original number is equal to 108